# Sparse Latent Factor Forecaster (SLFF) with Iterative Inference for Transparent Multi-Horizon Commodity Futures Prediction

### v4

@@ 3.3.1 Choosing $K$
We sweep $K\in\{1,\ldots,15\}$ on validation folds, monitoring $E$ and reconstruction errors. Figure 2 shows that energy decreases rapidly until $k\approx 8$ and flattens thereafter. We therefore use $K=10$ for the default setting and expose ablations in Section 5.4. The trend is consistent across commodities and seeds.
@@
3.4 Bridging Training and Test Time
During training, we compute $\mathbf{z}^{\star}_{t}$ using the iterative solver conditioned on the realized $Y_{t}$. At deployment, only $X_{t}$ is accessible; hence $\mathbf{z}_{t}=\text{Enc}_{X\rightarrow Z}(X_{t})$. To ensure that $\text{Enc}_{X\rightarrow Z}$ is not learning an impossible target, we:
  • minimize $\|\mathbf{z}^{\star}_{t}-\text{Enc}_{X\rightarrow Z}(X_{t})\|_{2}^{2}$ in alternating updates with the forecasting loss (two-stage loop, with a stop-gradient on $\mathbf{z}^{\star}_{t}$);
  • optionally (not in main experiments) add a contrastive regularizer that shuffles $Y_{t}$ within the batch while holding $X_{t}$ fixed, forcing the encoder-matched latents to be invariant to inconsistent futures; this prevents $\text{Enc}_{X\rightarrow Z}$ from chasing sample-specific $Y$ noise that $\mu$ alone would still allow via the refinement path;
  • include the proximity term $\mu\|\mathbf{z}-\text{Enc}_{X\rightarrow Z}(X_{t})\|_{2}^{2}$ in the inference objective and warm-start $\mathbf{z}^{(0)}=\text{Enc}_{X\rightarrow Z}(X_{t})$ to keep $\mathbf{z}^{\star}_{t}$ predictable from $X_{t}$;
  • treat $\text{Enc}_{X\rightarrow Z}(X_{t})$ as a stop-gradient inside the inference loop (default two-stage), so $\mathbf{z}^{\star}_{t}$ is computed with a frozen encoder snapshot and cached before encoder updates; this avoids a moving-target objective unless the unrolled end-to-end variant is explicitly enabled.
  • report alignment diagnostics (Table 10) to show that $R^{2}$ between $\mathbf{z}^{\star}_{t}$ and $\text{Enc}_{X\rightarrow Z}(X_{t})$ is around 0.73 on validation/test folds and that cosine similarity remains high on held-out splits.
Section 5.4 shows that removing this consistency training noticeably erodes both forecasting accuracy and interpretability.
3.5 Differentiable Variants
To address the concern that $\text{Enc}_{X\rightarrow Z}$ parameters were previously frozen, we now support three regimes:
  1. 1.
    Two-stage training (default, stop-grad inference): $\text{Enc}_{X\rightarrow Z}$ is treated as a stop-gradient inside the inference loop, so $\mathbf{z}^{\star}_{t}$ is computed with a fixed encoder snapshot. The encoder is then updated only via the match/contrast losses, not the forecast loss, which controls train-test mismatch and discourages a decoder that only works with optimization-refined latents.
  2. 2.
    Unrolled end-to-end (fully differentiable): We backpropagate through all $K$ inference steps and allow gradients to flow through $\text{Enc}_{X\rightarrow Z}$, akin to learned ISTA. This removes the stop-grad assumption and directly couples inference and prediction, but is $\sim 18\%$ slower; we include it in ablations for completeness.
  3. 3.
    Deployment fine-tune: after encoder alignment, we optionally fine-tune Pred and Dec using $\hat{\mathbf{z}}=\text{Enc}_{X\rightarrow Z}(X_{t})$ in the prediction loss to optimize the deployed path.
3.6 Training Objective and Algorithm
The total loss is
$L=L_{\text{pred}}+\beta L_{\text{match}}+\gamma L_{\text{contrast}},$ (6)
where $L_{\text{pred}}=C(Y_{t},\text{Dec}(\mathbf{z}^{\star}_{t},\mathbf{h}_{t}))$, $L_{\text{match}}=\|\mathbf{z}^{\star}_{t}-\text{Enc}_{X\rightarrow Z}(X_{t})\|_{2}^{2}$, and the contrastive term is defined by shuffling $Y_{t}$ within a batch while holding $X_{t}$ fixed:
@@ 3.6 Training Objective and Algorithm
which penalizes alignment between encoder latents and inconsistent future refinements. This complements the $\mu$ proximity constraint by discouraging a latent code that overfits to accidental $Y$ noise while remaining close to the encoder output. Unless noted, we set $\gamma=0$ in the main results, so the contrastive term is a diagnostic option and is not used in the main experiments. Hyperparameter $\beta$ is fixed at 5 across folds/commodities; $\gamma$ is set to 0 unless the contrastive regularizer is explicitly enabled. Unless Stage 3 is enabled, Pred/Dec are optimized on $\mathbf{z}^{\star}$ while evaluation uses $\text{Enc}_{X\rightarrow Z}(X)$; Table 13 in Appendix A reports the resulting diagnostic gap. We keep $L_{\text{pred}}$ on $\mathbf{z}^{\star}$ to preserve the refinement learning signal; mixed-loss or deployment-path fine-tune variants are optional and not used in the main tables. In the optional deployment fine-tune we replace $L_{\text{pred}}$ with $C(Y_{t},\text{Dec}(\hat{\mathbf{z}}_{t},\mathbf{h}_{t}))$ where $\hat{\mathbf{z}}_{t}=\text{Enc}_{X\rightarrow Z}(X_{t})$. Unless stated otherwise, reported results use the two-stage model without deployment fine-tuning (Stage 3 omitted).
The training loop proceeds as follows.
  1. 1.
    Sample a batch $\{(X^{(j)},Y^{(j)})\}$.
  2. 2.
    Compute $\mathbf{h}^{(j)}=\text{Pred}(X^{(j)})$.
  3. 3.
    Run $K$ inference steps (warm-started at $\text{Enc}_{X\rightarrow Z}(X^{(j)})$) to obtain $\mathbf{z}^{\star(j)}$, treating $\text{Enc}_{X\rightarrow Z}$ as stop-gradient unless the unrolled end-to-end variant is used.
  4. 4.
    Stage 1: update $\theta_{\text{Pred}},\theta_{\text{Dec}}$ using $L_{\text{pred}}$ (forecast loss with $\mathbf{z}^{\star}$).
  5. 5.
    Stage 2: cache $\mathbf{z}^{\star}$ computed from the frozen encoder snapshot and update $\theta_{X\rightarrow Z}$ using $L_{\text{match}}$ and $L_{\text{contrast}}$ against that fixed target.
  6. 6.
    Stage 3 (optional): fine-tune $\theta_{\text{Pred}},\theta_{\text{Dec}}$ using predictions from $\hat{\mathbf{z}}=\text{Enc}_{X\rightarrow Z}(X)$.
  7. 7.
    Early-stop based on validation RMSE aggregated across horizons and folds.
@@ 3.7 Interpretability via Structured Sparsity
We define an active factor as $|z_{k}|>\epsilon_{\text{active}}$ with $\epsilon_{\text{active}}=1\text{e-}3$, and report the mean count per timestamp. In the full-scale commodity experiments ($m=16$) this yields 5577 active factors, while the diagnostic $\mu$-sweep uses a reduced latent dimension ($m=3$) and therefore reports counts around 2233. To ensure factors are comparable across seeds we align $\mathbf{z}$ using an orthogonal Procrustes rotation and report stability scores (Section 3.8). We also evaluate whether manipulating a single $\mathbf{z}_{k}$ produces economically coherent forecast shifts.
@@ 3.8 Interpretability Protocol
We formalize interpretability as a three-stage protocol: (i) stability, measured by Procrustes-aligned subspace overlap across seeds and folds; (ii) driver validity, measured by out-of-sample regressions against observable indices not used by the decoder; and (iii) behavioral consistency, measured by counterfactual perturbations of $\mathbf{z}_{k}$ and event-study responses. This protocol is implemented in the released notebooks and summarized in Section 5.5 and Appendices CD.
@@ 4.1 Data Sources
4.2 Data Availability and Licensing
The full experimental dataset includes a mix of public and proprietary feeds. Public sources include EIA, BLS, FRED/ALFRED, IMF, World Bank, OECD, NOAA/NWS, CFTC, and exchange settlement data (CME/ICE/LME). Proprietary sources include Bloomberg, Refinitiv Eikon, S&P Global Market Intelligence, and GlobalData. Exact reproduction of all reported results therefore requires the corresponding licenses. To support reproducibility, we release (i) release-lag calendars, (ii) a machine-readable feature map with provenance, and (iii) a public-only pipeline that substitutes public proxies for proprietary series and provides stubs for licensed feeds. Appendix B lists the proxy mapping used in the public-only pipeline. The public-only pipeline reproduces the end-to-end methodology, but absolute error metrics may differ from the full dataset and are not claimed to match the headline results.
@@ 4.3 Information Set Definition and Leakage Mitigation
To guard against mixed-frequency leakage, we codify the “available information at end-of-day tt” using release calendars and vintage repositories. Table 3 summarizes representative feature blocks, their publication cadence, and the transformations used to align them with daily observations. Three principles are enforced across the pipeline:
@@
    Release-aware alignment: Each macro series is lagged by its actual publication delay (e.g., ISM PMI published on the first business day of the subsequent month enters the feature matrix on that date and is forward-filled only afterwards).
@@
    Vintage storage: We rely on ALFRED, FRED-MD vintages, IMF WEO historical releases, and EIA real-time STEO snapshots so that revisions are not backfilled retroactively.
@@
    Forward-fill discipline: Lower-frequency values are forward-filled only until the next scheduled release; the placeholder is masked for models to avoid signaling the future.
Table 3: Information set specification at the New York 17:$0$ ET cut-off.
Block Native Frequency Publication Lag Sources Alignment / Notes
Exchange prices, spreads, realized vol Daily / Intraday Same day CME/ICE/LME settlements, Bloomberg Refdata Cleaned end-of-day snapshots; realized vol uses only past prices up to $t$.
CFTC Commitment of Traders Weekly (Tue data, Fri release) +3 business days CFTC legacy and disaggregated reports Recorded on official release timestamp; values available only from release Friday onwards.
EIA petroleum status + inventories Weekly (Wed 10:30 ET) +2 business days EIA Weekly Petroleum Status Report API Feature timestamp is the official release date/time; global features are forward-filled until next report.
PMI/Industrial Production (ISM, Caixin, Eurostat) Monthly 1–6 weeks ISM, S&P Global, National Bureau of Statistics, Eurostat Each release is stored with ALFRED-style vintage IDs. Forward-fill begins only after the release date.
Inflation, employment, GDP revisions Monthly / Quarterly 2–8 weeks BLS CPI/PPI, BEA GDP, OECD, IMF WEO Final values never overwrite prior vintages; the dataset references the latest version known on date $t$.
Event indicators (OPEC+, geopolitical, strike tracker) Event-driven Real-time + manual coding Thomson Reuters Eikon, EIA alerts, GlobalData strike tracker Binary indicators and intensity scores recorded within 6 hours of publication; no hypothetical proxies.
Weather / climate indices (ENSO, hurricanes) Weekly / Monthly 0–2 weeks NOAA CPC, NHC Resampled to daily; release-lag aware fill.
For each feature block we generate an “availability mask” that flags whether the latest observation is stale relative to its scheduled release. The mask is concatenated to $X_{t}$, preventing the model from mistaking a constant forward-filled value for new information. The entire pipeline runs inside a reproducible Prefect flow, and release calendars are version-controlled for auditability.
4.4 Feature Selection
4.4.1 Target Variable (Daily Futures Prices)
Copper (LME nearby futures), Crude Oil (WTI, Brent front-month). We refer to WTI futures as WTI throughout for brevity.
4.4.2 Lagged Target Variables & Price Dynamics (Daily)
Historical prices, log returns, realized volatility [4], futures curve spreads [33].
4.4.3 Macroeconomic Indicators (Daily, Monthly, Quarterly)
Global/Major Economy PMIs, Industrial Production, inflation (CPI, PPI), interest/currency rates.
4.4.4 Supply/Demand Fundamentals (Varying Frequencies)
Copper: Exchange inventories, producer output, China consumption/imports, scrap indicators. Crude Oil: EIA weekly inventories, US/OPEC+ production, rig counts, demand estimates.
4.4.5 Market Sentiment & Financial Data (Daily, Weekly)
Futures volume/open interest, CFTC COT reports, VIX, OVX, S&P 500, FTSE 100, Shanghai Composite, MSCI EM currency baskets.
4.4.6 Technical Indicators (Calculated from Daily Prices)
SMAs, EMAs, RSI, MACD, ADX [66].
4.4.7 Other (Exploratory)
Inter-commodity spreads, weather indices (ENSO, hurricane impacts).
@@ 4.5 Data Frequency, Time Period, and Rolling Windows
All features are stored at the daily close. Mixed-frequency data (weekly EIA, monthly PMI, quarterly GDP) are mapped onto the daily grid using the release-aware procedure from Section 4.3. The raw dataset covers Jan 1, 2005–Dec 31, 2023, but the evaluation protocol begins in 2013 after a 7-year burn-in for lagged features and rolling statistics.
Table 4: Rolling-origin fold definitions (train: 6y, val: 1y, test: 1y).

### v5

@@ 3.3.1 Choosing $K$
We select $K=10$ based on validation sweeps showing energy convergence by $k\approx 8$ (Figure 2), balancing accuracy and compute.
@@
3.4 Optional: Contrastive Regularizer
A contrastive regularizer can optionally be applied during encoder training to prevent overfitting to sample-specific noise in $Y_{t}$. This term is disabled by default ($\gamma=0$) and included for completeness. When enabled, it shuffles $Y_{t}$ within the batch while holding $X_{t}$ fixed:
@@ 3.4 Optional: Contrastive Regularizer
which penalizes alignment between encoder latents and inconsistent future refinements. The full objective when enabled is $L=L_{\text{pred}}+\beta L_{\text{match}}+\gamma L_{\text{contrast}}$, with $\gamma=0$ in all main experiments. This optional extension and its ablation are reported in the supplement.
3.5 Bridging Training and Test Time
To align training (with $Y_{t}$) and test (without), we: (i) minimize $L_{\text{match}}$ in alternating updates, (ii) include proximity $\mu$ and warm-start, (iii) use stop-gradients on Enc in inference to avoid moving targets, (iv) report diagnostics (e.g., $R^{2}\approx 0.73$ in Table 11).
3.6 Differentiable Variants
  • Two-stage (default): Update Dec/Pred on $\mathbf{z}^{\star}$, then Enc on fixed $\mathbf{z}^{\star}$.
  • Unrolled: Backprop through $K$ steps.
  • Fine-tune: Optimize deployed path post-alignment.
3.7 Training Algorithm
Algorithm 3.7 specifies the full training procedure with explicit gradient-flow annotations.
Algorithm 1: SLFF Two-Stage Training Input: Dataset $\mathcal{D}$; hyperparameters $K,\alpha,\lambda,\mu,\beta$; optimizers $\text{opt}_{1}$ (for $\theta_{\text{Pred}},\theta_{\text{Dec}}$), $\text{opt}_{2}$ (for $\theta_{\text{Enc}}$) for epoch $1,\ldots,E_{\max}$ do
for each mini-batch $\{(X^{(j)},Y^{(j)})\}_{j=1}^{B}$ do
  // — Inference (stop-gradient on Enc throughout) —
  $\mathbf{h}^{(j)}\leftarrow\text{Pred}(X^{(j)};\theta_{\text{Pred}})$
$\bar{\mathbf{z}}^{(j)}\leftarrow\texttt{stop\_grad}(\text{Enc}(X^{(j)};\theta_{\text{Enc}}))$ // freeze Enc snapshot
  $\mathbf{z}^{(0)}\leftarrow\bar{\mathbf{z}}^{(j)}$ // warm-start
  for $k=0,\ldots,K-1$ do
   $g\leftarrow\nabla_{\mathbf{z}}\big[C(Y^{(j)},\text{Dec}(\mathbf{z}^{(k)},\mathbf{h}^{(j)}))+\mu\|\mathbf{z}^{(k)}-\bar{\mathbf{z}}^{(j)}\|_{2}^{2}\big]$
$\mathbf{z}^{(k+1)}\leftarrow\text{SoftThresh}_{\alpha\lambda}(\mathbf{z}^{(k)}-\alpha\,g)$ end for
  $\mathbf{z}^{\star(j)}\leftarrow\mathbf{z}^{(K)}$ // cache refined latent
  // — Stage 1: update Pred, Dec on refined latents —
  $L_{\text{pred}}\leftarrow\frac{1}{B}\sum_{j}C(Y^{(j)},\text{Dec}(\mathbf{z}^{\star(j)},\mathbf{h}^{(j)}))$
$\theta_{\text{Pred}},\theta_{\text{Dec}}\leftarrow\text{opt}_{1}.\text{step}(\nabla_{\theta_{\text{Pred}},\theta_{\text{Dec}}}L_{\text{pred}})$ // — Stage 2: update Enc to match cached $\mathbf{z}^{\star}$
  $L_{\text{match}}\leftarrow\frac{1}{B}\sum_{j}\|\texttt{stop\_grad}(\mathbf{z}^{\star(j)})-\text{Enc}(X^{(j)};\theta_{\text{Enc}})\|_{2}^{2}$
$\theta_{\text{Enc}}\leftarrow\text{opt}_{2}.\text{step}(\nabla_{\theta_{\text{Enc}}}\beta\,L_{\text{match}})$ end for
[…]
end for
Key implementation details: (i) stop_grad on Enc during inference means the proximity term $\mu\|\mathbf{z}-\bar{\mathbf{z}}\|_{2}^{2}$ uses a fixed encoder snapshot $\bar{\mathbf{z}}$ throughout all $K$ steps within each batch, so the quadratic anchor does not shift mid-loop. (ii) Stages 1 and 2 execute sequentially within each batch: Pred/Dec update first on the cached $\mathbf{z}^{\star}$, then Enc updates on the same cached $\mathbf{z}^{\star}$ (which is not recomputed after the Pred/Dec update). (iii) $\mathbf{z}^{\star}$ is refreshed every batch; there is no stale caching across batches. (iv) Optional Stage 3 (fine-tuning Pred/Dec on $\text{Enc}(X)$ outputs) is disabled in all reported experiments ($\gamma=0$).
@@ 3.8 Interpretability via Structured Sparsity
We define an active factor as $|z_{k}|>\epsilon_{\text{active}}$ with $\epsilon_{\text{active}}=1\text{e-}3$, and report the mean count per timestamp. In the full-scale commodity experiments ($m=16$) this yields 5577 active factors. To ensure factors are comparable across seeds we align $\mathbf{z}$ using an orthogonal Procrustes rotation and report stability scores (Section 3.9). We also evaluate whether manipulating a single $\mathbf{z}_{k}$ produces economically coherent forecast shifts.
@@ 3.9 Interpretability Protocol
We formalize interpretability as a three-stage protocol: (i) stability, measured by Procrustes-aligned subspace overlap across seeds and folds; (ii) driver validity, measured by out-of-sample regressions against observable indices not used by the decoder; and (iii) behavioral consistency, measured by counterfactual perturbations of $\mathbf{z}_{k}$ and event-study responses. This protocol is implemented in the released notebooks and summarized in Section 6.5 and Appendices AB.
@@ 4.1 Data Sources
Gold Price (benchmark) Monthly \$/oz LBMA (Prices) Freely available
@@ 4.2 Information Set Definition and Leakage Mitigation
To ensure no forward-looking bias in mixed-frequency data, we enforce:
@@
    Release-aware alignment: Lag each series by publication delay (e.g., PMI enters on release day, forward-filled thereafter).
@@
    Vintage storage: Use ALFRED/FRED-MD for revisions; no backfilling.
@@
    Forward-fill discipline: Fill only post-release; concatenate availability masks to $X_{t}$ signaling staleness.
Pipeline audited via Prefect flows with versioned calendars.
Worked example (three features on day $t$ = Jan 15, 2020):
  • Daily copper close: Available at end-of-day $t$; enters $X_{t}$ directly. No lag.
  • EIA weekly crude inventory (released Wednesdays for the prior week): Jan 15 is a Wednesday; the report released that morning covers data through Jan 10. Value enters $X_{t}$ on Jan 15 and is forward-filled through Jan 21 (next release). Availability mask bit: 1 on release day, 0 on forward-filled days.
  • ISM Manufacturing PMI (released 1st business day of month for prior month): December PMI was released Jan 2; that value persists in $X_{t}$ via forward-fill from Jan 2 through Jan 31. Mask bit: 1 on Jan 2, 0 on Jan 3–31. The January PMI is unavailable until Feb 3 and therefore never enters any $X_{t}$ before Feb 3.
Within the 60-day look-back window $X_{t}\in\mathbb{R}^{60\times d}$, each feature column thus contains the most recent released value on each historical day, with the binary mask column indicating recency. The decoder receives both the imputed values and the mask as input channels.
4.3 Feature Selection and Engineering
Exact d=67 features, selected via mutual information (>0.1>0.1 with targets) and low redundancy ($\rho<0.85$ within blocks), guided by domain priors:
  • Price Dynamics (18): Log returns, realized volatility (5, 20, 60-day windows), futures spreads (contango or backwardation), sources: CME and ICE settlements.
  • Macroeconomic (15): PMIs (ISM/Caixin), IP, CPI/PPI, rates (Fed/ECB), FX (DXY, CNY), sources: FRED, OECD.
  • Fundamentals (14): Inventories (LME/EIA), production (USGS/ICSG/OPEC), demand (DOE supplied, China imports), rig counts, sources: EIA, NBS.
  • Sentiment/Financial (10): COT positions, VIX/OVX, equity indices (S&P, SSE), EM currencies, sources: CFTC, Bloomberg.
  • Technicals (6): SMA/EMA (10/50/200), RSI (14), MACD, ADX, computed from prices.
  • Other (4): Inter-commodity spreads, ENSO/hurricane indices, sources: NOAA.
Gold adds 3 mining-specific (AISC, reserves, ETF flows from LBMA/WGC). Total shared: 64; per-commodity: +3.
@@ 4.4 Data Frequency, Time Period, and Rolling Windows
All features are stored at the daily close. Mixed-frequency data (weekly EIA, monthly PMI, quarterly GDP) are mapped onto the daily grid using the release-aware procedure from Section 4.2. The raw dataset covers Jan 1, 2005–Dec 31, 2025, but the evaluation protocol begins in 2013 after a burn-in for lagged features and rolling statistics.
Table 3: Rolling-origin fold definitions (train: 6y, val: 1y, test: 1y). Updated to 2025.